OpenStudy (anonymous):

Identify the boundaries of the solution of this inequality. |Y+9|>4

OpenStudy (anonymous):

because you are dealing with absolute value, consider when Y+9 is greater than 4, and when Y+9 is less than -4

OpenStudy (anonymous):

solve for Y+9 = 4 and Y+9 = -4, those will be the two points at which your inequality will not be satisfied, therefore any value greater than that found for Y+9=4 will be a solution, and any value less than that found for Y+9=-4 will be a solution, there you have the 2 boundaries